### v14

# Consequences of a Two-Time Relativistic Bohmian Model

Abstract
Effects of a Bohmian type quantum-relativistic theory are explored. The model is obtained by introducing a new and independent time parameter whose relative motions are not directly observable and cause the quantum uncertainties of the physical observables. Unlike the usual de Broglie-Bohm theories, the Quantum Potential does not affect directly the observable motion but it determines that one relative to the new time variable. It turns out that the Zitterbewegung of a free particle, of which a more general law is obtained, is the key example of these hidden motions and, through it, it seems possible to give physical reality to the Feynman’s paths. A relativistic revision of the uncertainty principle also derives from the theory.
Keywords: Quantum Mechanics Foundations; de Broglie-Bohm Theory; Zitterbewegung; Uncertainty principle verification; Extra dimensions.
mmm
I Introduction
At today, while the de Broglie-Bohm (or pilot-waves) theory [1, 2, 3, 4, 5, 6] enjoys some favor as a deterministic approach to non-relativistic Quantum Mechanics, the picture is not so favorable as regards the attempts at its relativistic generalization. For some time, the characteristics of the instantaneous action at a distance admitted by the theory has been considered a manifest violation of the Lorentz invariance [7, 8, 9, 10]. Indeed, it has been shown that this problem can be solved, but at the cost of admitting superluminality [11, 12, 13, 14, 15]. Currently, the most authoritative Lorentz-covariant pilot waves theories can be grouped in this way, in terms of the proposed strategy to define simultaneity [16]:
    Introducing a temporal parameterization that synchronizes the trajectories of all particles [11, 12, 13, 17, 18]. This approach has the drawback of not making concrete statistical predictions, although subsequent developments suggest this possibility under certain circumstances [15, 19, 20, 21].
    Defining a future light-cone as surface of simultaneity [22]. This theory is unable to predict any probability.
    Rejecting the decomposition of the wavefunction into norm and phase, and interpreting it as an real quantum field exerting force on the particle [23]. Here, tachyon particles are avoided but it is not clear if and how the model can be verified.
    Defining simultaneity through ”slicings” of the space-time (foliations) by spacelike hypersurfaces. The majority of researchers favor a preferred foliation, unique for all particles [24, 25, 26, 27, 28, 30] but there are also those who suggest a degenerate foliation [29] or an union of all possible foliations [31].
The theories of the last group are thought by some researchers to be the most promising way to obtain the standard statistical predictions [32, 33]. However, these models introduce additional spacetime structures and admit superluminal velocities (and hence travels backwards in time in some Lorentz frames) for massive particles, in violation of Special Relativity, although not of Lorentz covariance. A more general reason for dissatisfaction is the fact that all the mentioned theories only aim to deterministically reproduce the predictions of standard quantum mechanics; for an advantage, therefore, exclusively conceptual, superfluous from a strictly phenomenological point of view. On the other hand, there is no shortage of fundamental quantum topics, for example spin and Zitterbewegung [34, 35], waiting to be satisfactorily understood and explained. It would certainly be desirable to have a relativistic quantum theory that is able to propose, together with a deterministic ontology, predictions beyond the standard interpretation, possibly in full compliance with the spirit of Special Relativity.
In this article, after observing that the more spontaneous relativistic generalization of the original de Broglie-Bohm theory is unsatisfactory as regards the continuity of the Quantum Potential with the non-relativistic case, we explore the strategy of introducing a new independent time parameter as the only hidden variable, whose relative motions are recognized to be instantaneous with respect to the traditional time. In this deterministic and explicitly non-local approach, a more general Zitterbewegung, as motion in the new time variable caused by the Quantum Potential, arises; then, it would be not directly observable, in accordance with recent beliefs [36, 37, 38]. The model is potentially capable to assign full physical reality to the Feynman paths and predicts, as non-standard fact, a modification of the uncertainty principle: by depending on the direction of motion, it would become anisotropic. This property could be observed in high energy accelerators.
II Relativistic generalization issue for the quantum potential in case of constant momentum
In the de Broglie-Bohm non-relativistic theory [1, 2, 3, 4, 5, 6], the wavefunction:
$\psi(\vec{r},t)=R(\vec{r},t)\,e^{\frac{i}{\hbar}S(\vec{r},t)}$ (1)
with $R>0$ and S real functions, is inserted into the Schrödinger equation, yielding the following equations:
$\frac{\partial S}{\partial t}+\frac{(\vec{\nabla}S){}^{2}}{2m}\,+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}=0$ (2)
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}\frac{\vec{\nabla}S}{m})=0$ (3)
Identifying S with the Hamilton’s principal function, we have $\vec{\nabla}S=\vec{p}$ and $\frac{\partial S}{\partial t}=-H$, so eq. (2) means the Hamilton-Jacobi equation with Hamiltonian:
$H=\frac{p{}^{2}}{2m}+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}$ (4)
where the last addend represents an energy, called Quantum Potential, that takes into account the quantum effects:
$V_{Q}\equiv-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}$ (5)
According to the Bohm’s original proposal, the motion equation is:
$\frac{d\vec{p}}{dt}=-\vec{\nabla}V-\vec{\nabla}V_{Q}$ (6)
In the case of a free particle (V=0), both a solution with variable $V_{Q}$ and a stationary solution with positive constant $V_{Q}$ are admitted [39, 40].
The direct relativistic generalization of this approach, still for a free particle, pushes to write: $\vec{\nabla}S=\vec{p}=m\,\gamma\,\vec{v}$ and $\frac{\partial S}{\partial t}=-H=-m\,\gamma\,c^{2}-V_{Q}$, with $V_{Q}$ to be determined. We will start by assuming that $\vec{p}$ is constant, while for $V_{Q}$, and therefore H, a possible dependence on time.
From eq. (1), taking the gradient, then the divergence and dividing by $\psi$, we obtain:
$\frac{\nabla{}^{2}\psi}{\psi}=\frac{\nabla{}^{2}R}{R}+2\,\frac{i}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}-\frac{p^{2}}{\hbar{}^{2}}$ (7)
being $\vec{\nabla}\cdot\vec{v}=0$. Comparison with the Klein-Gordon equation:
$\nabla{}^{2}\psi-\frac{1}{c^{2}}\,\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{m{}^{2}c^{2}}{\hbar{}^{2}}\,\psi$ (8)
provides:
$\frac{\nabla{}^{2}R}{R}+2\,\frac{i}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}=\frac{m{}^{2}\gamma^{2}c^{2}}{\hbar{}^{2}}+\frac{1}{c^{2}\psi}\,\frac{\partial^{2}\psi}{\partial t^{2}}$ (9)
having applied the identity $p^{2}+m^{2}c^{2}=m^{2}\gamma^{2}c^{2}$.
On the other hand, deriving twice eq. (1) over time, one gets:
$\frac{1}{\psi}\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}-\frac{H^{2}}{\hbar{}^{2}}-\frac{i}{\hbar}\frac{\partial H}{\partial t}$ (10)
Substituting eq. (10) into (9) and matching real and imaginary parts, we have:
$H^{2}=m^{2}\gamma^{2}c^{4}+\hbar{}^{2}(\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\frac{\nabla{}^{2}R}{R})$ (11)
$c^{2}\frac{\vec{\nabla}R}{R}\cdot\vec{p}+\frac{H}{R}\,\frac{\partial R}{\partial t}+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (12)
Now we assume that the continuity equation (3), with $m\gamma$ instead of m, still holds. By employing $\vec{\nabla}\cdot\vec{v}=0$, we get just the expansion of $\frac{dR^{2}}{dt}$. Of course, we can refer it simply to R:
$\frac{\partial R}{\partial t}+\vec{v}\cdot\vec{\nabla}R=\frac{dR}{dt}=0$ (13)
as well as to any function of R. Indeed, since $\vec{v}$ is constant, eq. (13) also holds for any function of any-order derivative of $R$ over time or space, and thus for $V_{Q}$ and H. We are therefore assuming $\frac{dH}{dt}=0$, but with the possibility that $\frac{\partial H}{\partial t}\neq 0$, and the same for $V_{Q}$. Substituting eq. (13) into eq. (12) one finds:
$\frac{1}{R}\frac{\partial R}{\partial t}(H-m\gamma c^{2})+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (14)
where, looking for non-trivial solutions, one assumes $m\gamma\neq 0$, and $\frac{\partial R}{\partial t}\neq 0$.
The issue, now, is that equation of motion (6) implies a too drastic simplification of eq. (14). Indeed, it reads:
$\frac{d\vec{p}}{dt}=-\vec{\nabla}V_{Q}=0$ (15)
then, applying eq. (13) to $V_{Q}$, one obtains $\frac{\partial V_{Q}}{\partial t}=0$ and so it is also $\frac{\partial H}{\partial t}=0$. Therefore eq. (14) simply reduces to $H=m\,\gamma\,c^{2}$, finally resulting $V_{Q}=0$.
So, in the assumptions made, we have not general continuity for the Quantum Potential with the non-relativistic case.
III The two-time Bohmian model
In the first place it is therefore necessary to admit that, even in the case of a free particle, $\vec{p}$ is not constant. Indeed, inside the Heisenberg picture of the standard interpretation, Schrödinger and Dirac recognized that an omnidirectional velocity $\vec{v}_{i}$ of a sinusoidal motion of frequency $\frac{2m\gamma c{}^{2}}{\hbar}\,$ and amplitude about $\frac{\hbar}{2m\gamma c}\,$, known as Zitterbewegung [34, 35], must be superimposed on the actually observed velocity in average value. This motion has been much commented on, sometimes related to spin [41, 42, 43, 44] and Quantum Potential [45], but the debate on its exact interpretation, and even observability, does not seem yet resolved [36, 37, 38, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Hopefully, to rediscover this peculiar movement inside a deterministic theory could clarify its nature. It can only be due to the Quantum Potential; but equation of motion (6) still seems inadequate: in fact it has been shown that putting $\frac{d\vec{p}_{i}}{dt}=-\vec{\nabla}V_{Q}$ leads to inconsistency in [1 + 1] dimensions [40].
Here we explore the effect of introducing a new independent time variable for the spatial coordinates of the particle. Two time physics has been considered in various contexts. An additional temporal parameter is considered e.g. in [56, 57, 58, 59, 60] and, with a different approach, in [61, 62, 63, 64] in order to more accurately explain different quantum properties or obtain supersymmetries. But in these theories the new time is not independent from the old one as we wish.
Let’s admit, then, that the coordinates ($X,Y,Z$) of the particle depend not only on commonly experienced time t, but also on another, totally independent and hidden, time parameter $\tau$. Referring for simplicity only to $X(t,\tau)$, we get two velocities, one classic and one intrinsic:
$v_{cx}\equiv\frac{\partial X}{\partial t}\quad\quad v_{ix}\equiv\frac{\partial X}{\partial\tau}\,$ (16)
having so: $dX=v_{cx}dt+v_{ix}d\tau$. If we suppose that $v_{cx}$ and $v_{ix}$ depend, respectively, only on $t$ and $\tau$, we have:
[…]
having set $t_{0}=\tau_{0}=0$.
The intrinsic motion is not directly observable and, for the independence of $\tau$, is perceived as instantaneous with respect to time t. This explicitly non-local behavior has dramatic effects on the measurements of the observable quantities. As for the position, eq. (17) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this must be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to even more surprising facts, not limited to the cases of an electron spreaded in an atomic orbital, quantum diffraction or entanglement, but really in any motion. Indeed, according to Feynman’s sum over histories interpretation (which is equivalent to the standard one) the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled at once all the possible paths from A to B [65] [^1] Also interacting with every possible virtual particle of the vacuum, like experience excellently confirms..
For a particle in presence of a potential $V$ the wavefunction is then written:
$\psi(\vec{r},t,\tau)=R(\vec{r},t,\tau)\,e^{\frac{i}{\hbar}S(\vec{r},t,\tau)}$ (18)
The motion is decomposed into a classical and a purely quantum part. The law of classical motion is well known; in particular, if V is conservative, the classic total energy is constant:
$m\gamma_{c}c^{2}+V=const\equiv E_{c}$ (19)
and deriving over time one can get:
$\frac{d\vec{p}_{c}}{dt}=-\vec{\nabla}V$ (20)
where the Quantum Potential does not appear; that is also true, naturally, for its non-relativistic approximation. This equation is therefore incompatible with the Bohm’s motion equation (6). On the other hand, if we take the total derivative over $t$ of eq. (4), we re-obtain the eq. (20), because $H-V_{Q}=E_{c}=const$. The model is thus in disagreement with the traditional non-relativistic de Broglie-Bohm theory. However, due to its setting, we will continue to call it Bohmian.
For $t$-motion we also require the continuity equation:
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}{\vec{v_{c}})}=0$ (21)
As regards the omnidirectional intrinsic motions, it is spontaneous to consider that they are the direct cause of the quantum uncertainties. More, we assume that the Special Relativity does not hold for them. This condition is in agreement with the standard Zitterbewegung, where the mass of the particle does not undergo any relativistic increase - other than that one due to the classic velocity $\vec{v_{c}}$ despite it reaches the speed c. Being $m\gamma_{c}$ the mass involved in the generic intrinsic motion, we so have $\vec{p}_{i}=m\gamma_{c}\vec{v}_{i}$.
For the generic $\tau$-motion we also have to suppose the validity of:
$\frac{\partial R}{\partial\tau}-\vec{v_{i}}\cdot\vec{\nabla}R=0$ (22)
To justify it, we can observe that it is a form of continuity equation in which the addends subtract instead of adding up. This would remind the behavior of an antiparticle, although in the following we will not make any reference to antimatter and only use positive energy values. Since $\tau$-motion is not directly influenced by $V$, we can consider that $v_{i}$ only depends on $\tau$; consequently, eq. (22) holds for any function of any-order derivative of $R$ over $t$ or space. Hence we are admitting that it can be applied to $V_{Q}$, at least in case of conservative $V$.
Conservation of energy for a generic motion in $\tau$ reads:
$V_{Q}+\frac{1}{2}m\gamma_{c}v_{i}^{2}\equiv E_{q}$ (23)
representing a purely quantum total energy, independent explicitally on $\tau$. Taking the $\frac{\partial}{\partial\tau}$ of eq. (23) and using eq. (22) for $V_{Q}$, one gets:
$m\gamma_{c}\frac{d\vec{v}_{i}}{d\tau}=-\vec{\nabla}V_{Q}$ (24)
The classic motion equation (20) and the generic $\tau$-motion (24) have so to replace the Bohm’s original one (6).
The new Bohmian model can thus be so summarized:
  1. 1.
    The spatial coordinates of the particle have to vary as a function of two independent temporal parameters, t and $\tau$. Motions in $\tau$ occur in all directions; they are responsible for quantum uncertainties and not directly observable.
  2. 2.
[…]
  3. 3.
    For $t$-motion and a generic $\tau$-motion the continuity eqs. (21) and (22), respectively, hold.
  4. 4.
    The Bohm’s law of motion (6) is replaced by eqs. (20) and (24) to get, respectively, t-motion and generic $\tau$-motion; for the latter the Special Relativity is not valid.
Certainly, guidance equation $\vec{p}_{c}=\hbar\,Im\,(\frac{\vec{\nabla}\psi}{\psi})$ is obtained by taking the gradient of $\psi$ and applying $\vec{\nabla}S=\vec{p}_{c}$.
It should also be noted that, although each $\tau$-motion does not respect the Special Relativity, the maximum value of every $v_{i}$ is $c$: in fact it represents the quantum uncertainty on the homologous component of $v_{c}$, which cannot be greater than $c$.
The model therefore adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we will see in the next sections, the motions in the new time variable are absolutely negligible for macroscopic objects. Conversely, when this is not the case, it happens that, even assuming to know the Quantum Potential with a good approximation, due to the absolute independence of the two times and the omnidirectionality of the $\tau$-motion, the function $\vec{r}(t,\tau)$ is unable to provide us the actual trajectory followed by the particle: it rather represents a three-dimensional object of $t$-spacetime. In particular, while $t$ remains constant, the particle can $\tau$-oscillate an arbitrary number of times in any arbitrary direction. However, this is an indeterminism linked to the non-existence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them, but the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
IV Zitterbewegung in the new model
Remarkably, a Zitterbewegung motion in the $\tau$ variable emerges after three integrations by studying the free particle in a fixed direction.
Here, basically, we will follow the deduction made in [40]; but, thanks to the more general form of eq. (24), obtained using eq. (22), we will get a simpler and more rigorous derivation.
The classical motion equation (20) just informs us that $\vec{p}_{c}$ is constant. All equations found in section II - except (6) - are valid. Again, eq. (21) can be rewritten:
$\frac{\partial F}{\partial t}+\vec{v_{c}}\cdot\vec{\nabla}F=\frac{dF}{dt}=0$ (25)
being $F$ any function of any-order derivative of $R$ over $t$ or space. In particular it holds for $V_{Q}$ and $H$: they are therefore fields that travel in space, although with zero total derivative with respect to $t$-time.
We will begin by studying only the classical rectilinear trajectory, called $\vec{x}=x\hat{v_{c}}$. Here, eq. (25) becomes the equation of a progressive wave:
$\frac{\partial F}{\partial t}+{v_{c}}\dot{F}=0$ (26)
with a dot denoting the derivative with respect to x. Sure, deriving again, we obtain the d’Alembert equation:
$\frac{\partial^{2}F}{\partial t^{2}}-{v_{c}^{2}}\ddot{F}=0$ (27)
So, $V_{Q}$ and H are waves that follow the particle in $t$-motion. Reasoning in the same way on eq. (22), we find instead that they are regressive waves in $\tau$-motion.
Using the above equations for F=R, eqs. (11) and (14) can be rewritten as:
$H=\pm\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar{}^{2}\frac{c^{2}}{\gamma_{c}^{2}}\frac{\ddot{R}}{R}}$ (28)
$v_{c}\,\frac{\dot{R}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (29)
In the following we assume $m\gamma_{c}\neq 0$, $v_{c}\neq 0$, $\dot{R}\neq 0$ and limit to consider only non-negative values for the Hamiltonian. Substituting eq. (28) into eq. (29), we get:
$v_{c}\,\frac{\dot{R}}{R}\,(\sqrt{\beta}-\beta)=-\frac{1}{4}\frac{\partial\beta}{\partial t}$ (30)
where $\beta\equiv\frac{H^{2}}{m^{2}\gamma_{c}^{2}c^{4}}=1-\lambda^{2}\,\frac{\ddot{R}}{R}$, with $\lambda^{2}\equiv\frac{\hbar{}^{2}}{m^{2}c^{2}\gamma_{c}^{4}}$. Using eq. (26) for $F=\beta$, one finds:
$\frac{\dot{R\,}}{R}\,=\frac{1}{4}\,\frac{\dot{\beta}}{\sqrt{\beta}-\beta}$ (31)
Multiplying by dx and integrating member by member we can obtain:
$1+\frac{c_{1}}{R^{2}}=\sqrt{\beta}$ (32)
with $c_{1}$ arbitrary constant. So:
$H=m\gamma_{c}c^{2}\,(1+\frac{c_{1}}{R^{2}})$ (33)
and the Quantum Potential is $V_{Q}=m\gamma_{c}c^{2}\,\frac{c_{1}}{R^{2}}$. Compatibility with the stationary non-relativistic case requires $c_{1}>0$. By squaring eq. (32) one finds:
$(1+\frac{c_{1}}{R^{2}})^{2}=1-\lambda^{2}\,\frac{\ddot{R}}{R}$ (34)
Recalling that $R>0$, a first integration provides:
$\dot{R}=\pm\frac{1}{\lambda}\,\sqrt{\frac{c_{1}^{2}}{R^{2}}-4c_{1}ln\,R+c_{2}}$ (35)
with $c_{2}$ arbitrary constant. We look for a solution where R has a maximum, called $R_{M}$, at some point $x_{M}$, having in eq. (35) sign - for $x>x_{M}$ and + for $x<x_{M}$. From $\dot{R}(x_{M})=0$ we determine $c_{2}$ , getting:
$\dot{R}=\pm\frac{\sqrt{c_{1}}}{\lambda}\,\sqrt{\frac{c_{1}}{R^{2}}-\frac{c_{1}}{R_{M}^{2}}+4\,ln\,\frac{R_{M}}{R}}$ (36)
Placing $f\equiv\frac{c_{1}}{R_{M}^{2}}>0$, the Quantum Potential is rewritten:
$V_{Q}=m\gamma_{c}c^{2}f\,\frac{R_{M}^{2}}{R^{2}}$ (37)
and its minimum value is in $x_{M}$: $V_{Qm}=m\gamma_{c}c^{2}f$. R is determined by integrating eq. (36):
$\int\frac{dR}{\sqrt{f(\frac{R_{M}^{2}}{R^{2}}-1)+4\,ln\,\frac{R_{M}}{R}}}=-\frac{R_{M}\sqrt{f}}{\lambda}\mid x-x_{M}\mid+c_{3}$ (38)
which cannot be simplified by means of standard functions.
In a small neighborhood of the maximum, i.e. for $R\rightarrow R_{M}$, eq. (38) gives for R a quadratic dependence on $x-x_{M}$. Indeed, here the rooting at first member is approximated by $2(f+2)(1-\frac{R}{R_{M}})$, so by integrating we have:
$\sqrt{1-\frac{R}{R_{M}}}\simeq\frac{\sqrt{f(f+2)}}{\sqrt{2}\lambda}\,\mid x-x_{M}\mid+c_{4}$ (39)
where $c_{4}=0$ by imposing $R(x_{M})=R_{M}$. Hence:
$R\simeq R_{M}(1-\frac{f(f+2)}{2\lambda^{2}}(x-x_{M})^{2})$ (40)
$\frac{1}{R^{2}}\simeq\frac{1}{R_{M}^{2}}\frac{1}{1-\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2}}\simeq$
$\simeq\frac{1}{R_{M}^{2}}(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2})$ (41)
Then, from eq. (37) the Quantum Potential around $x_{M}$ can be approximated by the potential of a harmonic oscillator:
$V_{Q}\simeq m\gamma_{c}c^{2}f\,(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2})$ (42)
By replacing it into the $\tau$-motion equation (24) and imposing $x_{M}=X(t,0)=X_{t_{0}}+X_{\tau_{0}}+v_{c}\,t$, we ultimately get the following solutions:
$X(t,\tau)\simeq X_{t_{0}}+X_{\tau_{0}}+v_{c}\,t+\Delta X\,sin(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (43)
$v_{ix}(\tau)\simeq\Delta X\,\frac{c}{\lambda}\,f\sqrt{2(f+2)})\,cos(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (44)
where, according to point 1 of the theory, the amplitude $\Delta X$ represents the positional indeterminacy in $x$ of the particle. To get amplitude and angular frequency of the standard Zitterbewegung, i.e. $\frac{\hbar}{2m\gamma_{c}c}$ and $\frac{2m\gamma_{c}c{}^{2}}{\hbar}$, it is enough to set $\gamma_{c}\,f\sqrt{2(f+2)}=2$, and $v_{ix_{MAX}}=c$. However, by imposing only $v_{ix_{MAX}}=c$, we obtain $\Delta X=\frac{\lambda}{f\sqrt{2(f+2)}}$; as for f, we will simply assume that it is of the order of unity. Then, in the model under consideration, the uncertainty principle is:
$\frac{\lambda}{f\sqrt{2(f+2)}}\times m\gamma_{c}c=\frac{\hbar}{f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}}$ (45)
with an unexpected Lorentz factor to divide. Looking at eq. (28) this seems at all correct, because there is not only a $\gamma_{c}$ which increases the relativistic mass but also another $\gamma_{c}$ which decreases the influence of $\hbar$. At ultrarelativistic speeds, the quantum positional uncertainty in direction of motion should be negligible.
The amplitude of the $\tau$-oscillation - i.e. the quantum positional uncertainty - is inversely proportional to the rest mass. Then, for a particle of one picogram we find it about $15$ orders of magnitude smaller than for an electron. However, this is not the only reason that explains the evanescence of the $\tau$-motion for macroscopic bodies, as we will show in the next section.
The considered $\Delta X$ represents the minimum positional uncertainty of the particle; only by means of a wave packet description is it possible to take into account larger positional uncertainties (and smaller momentum uncertainties).
For $x=X$ in eq. (42), we get the dependence of $V_{Q}$ on $\tau$: it oscillates, and so does $H$, with an amplitude $\frac{1}{2}m\gamma_{c}c^{2}$; this value represents so the quantum uncertainty in a measure of energy, for a free particle with minimal positional uncertainty. Of course, multiplying by half-period of oscillation - which represents the quantum uncertainty in a measure of $t$-time - we find again the uncertainty principle:
$\frac{1}{2}m\gamma_{c}c^{2}\times\frac{\pi\lambda}{f\sqrt{2(f+2)\,}c}=\frac{\pi\hbar}{2\,f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}}$ (46)
The total purely quantum energy $E_{q}\equiv V_{Q}+\frac{1}{2}m\gamma_{c}v_{ix}^{2}$ is independent of $x$:
$E_{q}=V_{Qm}+\frac{1}{2}m\gamma_{c}c{}^{2}=m\gamma_{c}c^{2}(f+\frac{1}{2})$ (47)
and it also is the maximum value of $V_{Q}$, corresponding to the minimum value of R: $R_{m}=\frac{R_{M}}{\sqrt{1+\frac{1}{2f}}}$. Certainly, a normalization of $R{}^{2}$ (not exact, for the non-integrability of eq. (38)) could determine $R_{M}$.
As an important result it was therefore found not simply that the Quantum Potential is never negligible, but that its maximum value, which represents the total purely quantum energy, is at least of the order of $m\gamma_{c}c^{2}$, regardless of the value of $f$. Nonetheless, the $\tau$-motions can be negligible in amplitude, as happens for a macroscopic body. It is also surprising that, although the existence of $V_{Q}$ is due to $\hbar$, its maximum and minimum values do not depend on it: only the frequency of its oscillation does.
[…]
IV.1 Zitterbewegung in arbitrary direction
In an arbitrary direction $\hat{s}$, which forms an angle $\vartheta$ with $\hat{v}_{c}$, eq. (25), for F=R, is written:
$\frac{\partial R}{\partial t}+{v_{cs}}\dot{R}=0$ (48)
where $v_{cs}=v_{c}\,cos\,\vartheta$ and the dot denots derivative over s. So eqs. (28) and (29) are generalized by:
$H=\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar{}^{2}\frac{c^{2}}{\gamma_{cs}^{2}}\frac{\ddot{R}}{R}}$ (49)
$v_{cs}\,\frac{\dot{R}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (50)
Placing: $\lambda^{2}\equiv\frac{\hbar{}^{2}}{m^{2}c^{2}\gamma_{c}^{2}\gamma_{cs}^{2}}$, we get back the same equations we have considered. Expressions of the Hamiltonian and Quantum Potential, given by (33) and (37), do not change, not even the amplitude of their oscillations. By imposing the motion equation:
$m\gamma_{c}\frac{d{v}_{is}}{d\tau}=-\dot{V}_{Q}$ (51)
we obtain an intrinsic harmonic motion described again by (43) and (44), but with the new value of $\lambda$. The consequence on the uncertainty principle - eqs. (45, 46) - is the presence of $\gamma_{cs}$ instead of $\gamma_{c}$ to divide. This gives rise to a spatial anisotropy of the uncertainty principle:
$\Delta s\cdot\Delta p_{s}\sim\frac{\hbar}{\gamma_{cs}}$ (52)
In particular, for $\hat{s}$ tending to be perpendicular to $\hat{v}_{c}$, $\gamma_{cs}\rightarrow 1$ and the standard uncertainty principle, independent of speed, is re-established. However, the case of exact perpendicularity is singular: for $v_{cs}=0$ eq. (50) implies $H=const\Rightarrow V_{Q}=const\Rightarrow\vec{v}_{is}=const$, instead of being oscillatory.
IV.2 Non-relativistic limit. Singularity in $v_{c}=0$
For the non-relativistic ($v_{c}\ll c$) and $v_{c}\rightarrow 0$ cases, it is sufficient to approximate $\gamma_{c}$ in the previous equations (respectively, with $1+\frac{v_{c}^{2}}{2c^{2}}$ and 1), obtaining still high frequency oscillatory motions in $\tau$. These results cannot be obtained starting from the standard Schrödinger equation because from it the non-relativistic approximation of eq. (50) cannot be deduced. This happens since the explicit dependence of $R$ on $t$ causes the Klein-Gordon equation to no longer tend to the Schrödinger equation.
In $v_{c}=0$ there is another singularity; in fact, by eq. (50) we get again $\vec{v}_{is}=const$, contrary to the case $v_{c}\rightarrow 0$.
IV.3 Standard Zitterbewegung
Interpretating the oscillatory motion in $\tau$ as occurring in t, the standard Zitterbewegung is re-obtained. In the standard view, $R$ is considered time independent, so the second side of eq. (10) is just $-\frac{H^{2}}{\hbar{}^{2}}$, with $H=const$. In the new theory this means demanding: $\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}=0$, that provides $R\propto\,e^{\frac{2i}{\hbar}H\,t}$ and hence $V_{Q}\propto e^{-\frac{4i}{\hbar}H\,t}$. Then, using eqs. (22) for $V_{Q}$ and (24) (both as for t-time), after having conveniently determined the arbitrary constants, one can get: $v_{i}\propto c\,e^{-\frac{2i}{\hbar}H\,t}$, as found in the Heisenberg picture.
V Working account for the double-slit experiment
The described model, without questioning the traditional conservation laws, also adopts the standard evolution of the particle wavefunction through the Klein-Gordon equation. The introduction of the independent time parameter, which is hidden, basically provides a deterministic explanation of the uncertainty principle. We have seen that the latter changes in the direction of motion only at relativistic speeds. Excluding for now this case, which will be discussed in the next section, it can be stated that the empirical predictions of standard Quantum Mechanics are fully respected. For example, no difference is foreseen for the kinetic theory of gases, whose standard quantum model is obtained by discretizing the phase space into volumes $h^{\emph{3}}$. However, the model is able to resolve the paradoxical dualism that afflicts the standard interpretation, by explaining deterministically the collapse of the wave packet after a measurement. We will now detail how the theory works in the case of the most famous quantum experience: the double-slit experiment.
The standard wave packet description, compliant the de Broglie relation, is still empirically justified on the basis of the physical indeterminations of the sent particle. At the initial instant, we assume for simplicity that the particle has the minimum positional uncertainty previously found, which we call $\Delta s_{0}$. During motion, as long as the energy of the particle does not change, the growth of the positional uncertainty, obtainable by the Liouville’s theorem, continues to be the standard one: $\Delta s=\sqrt{\Delta s_{0}^{2}+\frac{\Delta p_{s0}^{2}}{m^{2}}\,t^{2}}$. At a fixed instant $\bar{t}$, $\Delta s$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-motions, is spread; in other words, at instant $\bar{t}$ it is simultaneously present in every point of this volume. We highlight that:
  1. 1.
    This simultaneity is not in conflict with Special Relativity, because the generic $\tau$-motion itself, as long as the energy of the particle does not change, does not constitute an event: $\bar{t}$ is fixed in any reference system.
  2. 2.
    Nothing prevents these infinite self-particles from interfering with each other in all possible physical ways. Mutual exchanges of particles correspond to what in the standard point of view is described as emission and reabsorption of virtual particles, obtaining excellent experimental verifications.
  3. 3.
    In addition to interacting like particles, self-particles interfere like waves, exactly as predicted by the standard interpretation, according to the Feynman interpretation [67]. The model simply specifies that the different histories are realized in time $\tau$. This description is in conflict neither with the interactions just discussed nor with the corpuscular description given by equations (20) and (24). Due to the constitutional limitations of the latter, the effective trajectory followed by the particle is everywhere unknowable and the wave description is necessary. When the particle reaches the wall, diffraction is observed if the sphere-like volume includes the two slits (Fig. 1).
Figure 1: Interpretation of the two slits experience according to the Two-Time Model: with respect to the hidden time $\tau$, independent of the normal time $t$, the particle reaches the slits at two different instants; but with respect to $t$ its presence in them is simultaneous.
Note that zig-zag trajectories which violate the conservation of $t$-momentum and yet admitted as legitimate Feynman paths, can now be properly physical, being able to be obtained by composition of $\vec{p}_{c}$ with one or more $\vec{p}_{is}$.
With respect to time $t$ (that is, relative to an observer), the particle travels at once all the possible paths between the initial point and the slits, provided that they are included in the volume swept by the sphere-like figure. In this, therefore, the model differs from the standard sum of histories interpretation, where there is no limitation for the paths.
V.1 Collapses
Areas where the self-interference are destructive constitute ”shadow zones” of the sphere-like volume, where from a corpuscular point of view, the Quantum Potential pushes the self-particles elsewhere. Dually, there will also be areas where the interference is constructive.
The apparent observational ”collapse” is an extreme case of destructive interference that occurs after a energy variation of the particle for collision (more generally, interaction) with an external particle (or field): it happens that the self-wavefunctions in the impact area change their phase; then, their interference with the self-wavefunctions located elsewhere cancels - instantaneously, with respect to $t$-time - the latter. From a corpuscular point of view, the Quantum Potential no longer allows the particle to $\tau$-move away from the proximity of the impact.
So, a non-zero energetic interaction with an external particle or field constitutes an observable (but not necessarily observed) event, capable of dramatically restricting $\Delta s$, its minimum value being of the Compton length order. From a macroscopic view, the intrinsic movements stop. Observation itself has no special role: this happens regardless of whether the external particle is from a detector or residual air.
Consistent with this description, after a perfectly elastic collision with a massive wall (it is also the case of a photon hitting a perfectly reflective mirror) the particle does not reduce its delocalization: no energy is transferred to the wall and the $\tau$-motions amplitude does not collapse.
This picture also indicates another fundamental reason that generally prevents a macroscopic body from spread out in space through intrinsic motions: compared to a particle, it is vastly more likely that it interacts energetically with other matter or fields. Indeed, even a very small variation in the phase of the wave is capable of making it collapse in the way described, canceling out significant quantum uncertainties.
For what has been said, the paradox that the particle, being $t$-simultaneously in every point of a certain macroscopic volume, violates the conservation of mass and of number of particles is only apparent: a particle search experiment will always detect only one particle. The total purely quantum energy distributed in the sphere-like volume in a $\Delta\tau$ interval is indeed infinite, but it is beyond any kind of experimentation. Really it is virtual, in the same sense used in standard view. Indeed, the existence of the hidden time $\tau$ is capable of explaining why the increasingly larger energies predicted by entering the infinitesimal are real (as experiments excellently confirm) but never directly observed.
VI Discussion
In Feynman’s (however standard) formulation of Quantum Mechanics, when a particle moves from one point to another, it is as if it followed at once all possible paths between the two points. In this article we have begun to explore a model in which, by abandoning the ”as if”, the particle really does so. For this to be possible without contradiction, it is necessary to add a new time parameter, totally independent of the commonly experienced one [^2] After all, time could logically be defined as the parameter that allows a particle to be at different points.. Called it $\tau$, the wavefunction of the particle is represented by eq. (18). In it, the interpretation of S does not change, so all equations referring to time t, as they use partial derivatives, consist just of a relativistic generalization of the classical Bohmian ones. Conversely, by assuming for the generic $\tau$-motion an appropriate continuity equation, new laws of motion, which must replace that one of Bohm’s original theory, derive. The $\tau$-motions are governed by the Quantum Potential and spread out the particle in a region of space that depends on the conditions of the physical surrounding. For a free particle, a more general Zitterbewegung motion has been found, having minimum amplitude of the order of the Compton length. In presence of a $V\neq 0$, the Quantum Potential is modified through eqs. (11) and (12) and the spreading through eq. (24); in the case of an atomic electron, for example, the $\tau$-motions justify the instantaneous dispersion of the electron in the orbit (we intend to show this in a subsequent publication).
In the model, the quantum uncertainties and the uncertainty principle itself have a deterministic explanation: they are due to the $\tau$-movements. For the latter, the Special Relativity does not hold, but since they are judged as instantaneous with respect to t-time, measured in any reference system, it is impossible to detect violations of the Lorentz covariance.
We have to emphasize that, since $m\gamma_{c}\neq 0$ is the only hypothesis made about the particle, all the conclusions obtained - including the characteristics of the oscillatory motions in $\tau$ continue to be valid for photons (for which, on the other hand, quantum diffraction figures analogous to those of any other particle are found).
The imposition of eq. (22), necessary to deduce the general equation (24), might seem a weakness of the theory. Anyway, it is legitimate that even the generic movement in $\tau$ possesses own equation of continuity, like that one in t. Its peculiarity, to describe a regressive wave, reminds the motion of an antiparticle.
A question concerns the determination of the parameter f. By assuming f = 1/2, one gets, in any direction, a total purely quantum energy $m\gamma_{c}c^{2}$, which could represent a virtual antiparticle. But, naturally, the determination of f should derive by rigorous considerations and it is not excluded that it may depend on the type of particle and/or its properties here not considered.
The evanescence of the uncertainty principle in direction of motion for $v_{c}\rightarrow c$, is recognizable just by observing the general eq. (49): there, a Lorentz factor - equal to 1 only in direction tending to be perpendicular to the rectilinear classic motion - directly reduces the influence of $\hbar$. Then, the spreading region of the particle should be contracted in direction of motion. Believing that the high energy available in today’s accelerators should have already detected deviations based on eq. (52), would be quite hasty. It is necessary to consider not only that high speeds make localization measurements hard but above all that simultaneous measurements of complementary observables are prohibited by the same principle of causality in the Special Relativity. These and several other problems make it very difficult to experimentally verify the uncertainty principle [69, 68].
Compared to spacetime structures planned by the current most accredited theories, introduction of a second time variable is a more radical but also mathematically simpler criterion which, neither logically nor physically, differ in essence from the introduction of an extra spatial dimension. In addition to the explanation of the Zitterbewegung, it predicts new physical effects, permitting, inside the Feynman’s formulation, to clarify the phase of the paths (redefinable as real physical trajectories, although unobservable) and hopefully to represent the sum over histories as a properly mathematical integral.
Conclusively, in the explored theory the apparently paradoxical aspects of Quantum Mechanics are just the effect of the existence of an extra unobservable time dimension. By traveling in a hidden extra-time dimension, you can deterministically explain why a particle or photon is able to interact instantly with more slits in a diffraction, to run across multiple paths between two points, to admit entanglements with other arbitrarily distant particles… All observed facts.